# Learning to Learn for Few-shot Continual Active Learning

### v2

@@ Abstract
Continual learning strives to ensure stability in solving previously seen tasks while demonstrating plasticity in a novel domain. Recent advances in CL are mostly confined to a supervised learning setting, especially in NLP domain. In this work, we consider a few-shot continual active learning (CAL) setting where labeled data are inadequate, and unlabeled data are abundant but with a limited annotation budget. We propose a simple but efficient method, called Meta-Continual Active Learning. Specifically, we employ meta-learning and experience replay to address inter-task confusion and catastrophic forgetting. We further incorporate textual augmentations to ensure generalization. We conduct extensive experiments on benchmark text classification datasets to validate the effectiveness of the proposed method and analyze the effect of different active learning strategies in few-shot CAL setting. Our experimental results demonstrate that random sampling is the best default strategy for active learning and memory sample selection to solve few-shot CAL problems.
@@
continual learning, meta-learning, active learning, text classification
@@ I Introduction
Continual learning (CL) aims to address the stability-plasticity dilemma in the context of sequential learning. Stability refers to the ability to alleviate the decline in model performance on previously learned tasks, in other words, preventing catastrophic forgetting [1]. Notably, inter-task confusion [2] is one of the major reasons that cause catastrophic forgetting. It is a phenomenon where the model confuses classes from different tasks. Plasticity refers to the ability to adapt to a new task. However, most existing CL methods assume that all data are labeled and sufficient for learning each task, which is impractical.
In real-world scenarios, labeled data are scarce and unlabeled data are abundant. The annotation cost, especially for NLP tasks, is often expensive. To narrow this gap, we consider a more realistic setting for CL, where we provide a small amount of labeled data and a large amount of unlabeled data with limited annotation budget for each CL task. We formulate this problem as a few-shot continual active learning (CAL) process.
@@
  • To the best of our knowledge, our paper is the first to address few-shot continual active learning in the NLP domain. The data stream is arbitrary and inherently unpredictable. The label space of a forthcoming task may or may not overlap with that of previous tasks. Hence, we formulate a few-shot continual active learning setting to consider both class-incremental learning (i.e., disjoint label space) and domain-incremental learning (i.e., overlapped label space) in one data sequence.
@@ III Preliminaries
In this work, we focus on an online task-free class-incremental learning scenario, in which training data stream passes only once without the presence of “task boundaries” [17]. Based on the task-free setting, we formulate the problem of few-shot continual active learning as follows. Assume that training stream consists of $K$ tasks, $\{\mathcal{T}_{1},\mathcal{T}_{2},...,\mathcal{T}_{t},...,\mathcal{T}_{T}\}$. Each task $\mathcal{T}_{t}$ has a small set of labeled instances $\mathcal{D}^{label}_{t}=\{(x_{i},y_{i})\}_{i=1}^{N_{t}\times K}$ and a pool of unlabeled data $\mathcal{D}^{pool}_{t}=\{(u_{i})\}_{i=1}^{|\mathcal{D}^{pool}_{t}|}$. The labeled set consists of $N_{t}$ classes, each with $K$ samples.
@@ III-A Label Space
Based on the label space $\mathcal{Y}$ of tasks, typical CL scenarios have domain-incremental learning, where $\mathcal{Y}_{t}=\bigcup_{i=1}^{T}\mathcal{Y}_{i}$ and class-incremental learning, where $\mathcal{Y}_{t}\cap\mathcal{Y}_{t}^{\prime}=\emptyset,\forall t\neq t^{\prime}$. Since the data stream is arbitrary, the label space $\mathcal{Y}_{k}$ of a task may or may not be disjoint from the other tasks. Hence, we consider both domain-incremental learning and class-incremental learning in one sequence.
@@ III-E Few-shot CAL Challenges
Few-shot CAL problems pose challenges in optimizing neural networks. The main challenges are: (a) quickly adapting the model to the current task $\mathcal{T}_{k}$ while alleviating catastrophic forgetting; (b) addressing inter-task confusion with data scarcity; (c) exploiting efficient active learning strategy that makes the best use of available data to solve continual learning problems.
@@ V-B Training
Each epoch has $m$ batches of examples instantaneously drawn from data stream. For each task, our model is trained on $\mathcal{D}^{label}$ first, and then on $\mathcal{D}^{pool}$. MAML consists of two optimization loops:
@@ VI-B Implementation details
[…]
@@ VI-D Evaluation Metrics
Figure 1: T-SNE visualization of memory samples in the hidden representation space. Data points with darker colors (purple, violet and pink) represent samples from earlier tasks. The black-circled data points belong to the last task. We provide accuracy for a better comparison.
@@ VII Further Analysis
Figure 2: Per task accuracy at different learning stages. Darker color indicates higher accuracy. From left to right, the color for each task progressively fades away, indicating forgetting while learning more tasks. UNC shows a lighter color compared to other AL methods, indicating a higher degree of forgetting. The red-outlined box in (a) shows the accuracy on AGNews (Task 2) after learning DBpedia (Task 3).
@@ VIII Conclusion
This paper considers a realistic scenario, namely few-shot CAL. We devise a novel method, i.e., Meta-Continual Active learning, to address the problem of few-shot CAL. Our model dynamically selects worthwhile unlabeled data to annotate, thereby preventing catastrophic forgetting and maximizing model performance. Specifically, the proposed model employs MAML with experience replay to improve stability and plasticity. We use textual augmentations to enhance intra- and inter-task generalization. In experiments, we examine the proposed model using different active learning strategies. The experimental results demonstrate the effectiveness of our method to solve few-shot continual active learning problems. However, the annotation budget is allocated equally for each task. The annotation budget allocation strategies can be further studied. Furthermore, we can extend our setting to other NLP tasks as future works.

### v3

@@ Abstract
Continual learning strives to ensure stability in solving previously seen tasks while demonstrating plasticity in a novel domain. Recent advances in CL are mostly confined to a supervised learning setting, especially in NLP domain. In this work, we consider a few-shot continual active learning (CAL) setting where labeled data are inadequate, and unlabeled data are abundant but with a limited annotation budget. We propose a simple but efficient method, called Meta-Continual Active Learning. This method sequentially selects the most informative examples from a pool of unlabeled data and requests labels to enhance performance. Specifically, we employ meta-learning and experience replay to address inter-task confusion and catastrophic forgetting. We further incorporate textual augmentations to ensure generalization. We conduct extensive experiments on benchmark text classification datasets to validate the effectiveness of the proposed method and analyze the effect of different active learning strategies in few-shot CAL setting. Our experimental results demonstrate that random sampling is the best default strategy for active learning and memory sample selection to solve few-shot CAL problems.
@@
continual learning, meta-learning, few-shot learning, active learning, text classification
@@ I Introduction
Continual learning (CL) aims to address the stability-plasticity dilemma in the context of sequential learning. Stability refers to the ability to alleviate the decline in model performance on previously learned tasks, in other words, preventing catastrophic forgetting [1]. Notably, inter-task confusion [2] is one of the major reasons that cause catastrophic forgetting. It is a phenomenon where the model confuses classes from different tasks. Plasticity refers to the ability to adapt to a new task. However, the majority of existing CL methods presume the availability of labeled data in abundance, deeming it adequate for learning every task. Their performance heavily relies on the large quantity and high quality of labeled data, which is impractical.
In real-world scenarios, labeled data are scarce and unlabeled data are abundant. The cost of annotation, particularly concerning NLP tasks, tends to be prohibitively high. To narrow this gap, we contemplate a setting for continual learning that aligns more closely with real-world scenarios. Herein, we provide a small amount of labeled data and a large amount of unlabeled data with limited annotation budget for each continual learning task. We formulate this problem as a few-shot Continual Active Learning (CAL) process. Concretely, model sequentially selects the most worthwhile data from a pool of unlabeled data and requests labels to enhance performance and solve continual learning problems.
@@
  • Our work tackles few-shot continual active learning challenges in the NLP domain. Since the data stream is arbitrary and inherently unpredictable, the label space of a forthcoming task may or may not overlap with that of previous tasks. Hence, we formulate a task-free few-shot continual active learning setting to consider both class-incremental learning (i.e., disjoint label space) and domain-incremental learning (i.e., overlapped label space) in one data sequence.
@@ III Preliminaries
In this work, we focus on an online task-free class-incremental learning scenario, in which training data stream passes only once without the presence of “task boundaries” [17]. Based on the task-free setting, we formulate the problem of few-shot continual active learning as follows. Assume that training stream consists of $T$ tasks, $\{\mathcal{T}_{1},\mathcal{T}_{2},...,\mathcal{T}_{t},...,\mathcal{T}_{T}\}$. Each task $\mathcal{T}_{t}$ has a small set of labeled instances $\mathcal{D}^{label}_{t}=\{(x_{i},y_{i})\}_{i=1}^{N_{t}\times K}$ and a pool of unlabeled data $\mathcal{D}^{pool}_{t}=\{(u_{i})\}_{i=1}^{|\mathcal{D}^{pool}_{t}|}$. The labeled set consists of $N_{t}$ classes, each with $K$ samples.
@@ III-A Label Space
Based on the label space $\mathcal{Y}$ of tasks, typical continual learning scenarios have domain-incremental learning, where $\mathcal{Y}_{t}=\mathcal{Y}_{t}^{\prime},\forall t\neq t^{\prime}$ and class-incremental learning, where $\mathcal{Y}_{t}\cap\mathcal{Y}_{t}^{\prime}=\emptyset,\forall t\neq t^{\prime}$. Due to the unforeseen nature of sequential learning, the label space $\mathcal{Y}_{t}$ of a task may or may not be disjoint from the other tasks. Hence, we allow $\mathcal{Y}_{t}\cap\mathcal{Y}_{t}^{\prime}\neq\emptyset$ or $\mathcal{Y}_{t}\cap\mathcal{Y}_{t}^{\prime}=\emptyset,\forall t\neq t^{\prime}$ to occur.
@@ III-E Few-shot CAL Challenges
The few-shot CAL setting poses hurdles for the optimization of neural networks. The main challenges are:
  • Quickly adapting the model to the current task while alleviating catastrophic forgetting, with limited labeled examples;
  • Addressing inter-task confusion with data scarcity;
  • Exploiting an efficient active learning strategy that makes the best use of available data to solve continual learning problems.
@@ V-B Training
Each episode has $m$ batches of examples instantaneously drawn from data stream. For each task, our model is trained on $\mathcal{D}^{label}$ first, and then on $\mathcal{D}^{pool}$. MAML consists of two optimization loops:
@@ VI-B Implementation details
Our example encoder is a pretrained BERTBASE model [28]. The parameter size of our model is 109M. For learning rates, $\alpha=1e^{-3}$ and $\beta=3e^{-5}$. The training batch size is 16. The number of mini-batches in each episode, $m$ = 5. The label budget $B_{A}$ = 2000 examples per task and the memory budget $B_{\mathcal{M}}$ is 5 per class, i.e., $n_{s}=5$. For textual augmentation, we swap words randomly as the weak augmentation. And, we apply the combination of swapping words randomly, deleting words randomly and substituting words by WordNet’s synonym, as the strong augmentation. We use nplug [^1] https://github.com/makcedward/nlpaug 1.1.10, a Python package to implement augmentations. All models are executed on Linux platform with 8 Nvidia Tesla A100 GPU and 40 GB of RAM. All experiments are performed using PyTorch [^2] https://pytorch.org/ 1.10.0+cu113 [29].
@@ VI-D Evaluation Metrics
Figure 1: T-SNE visualization of memory samples with different AL strategies. A equal dispersion of the data indicates a good memory representation. We provide accuracy for a better comparison. Data points with darker colors (purple, violet and pink) represent samples from earlier tasks. The black-circled data points belong to the last task. The clustering of memory samples from the last task suggests model focuses more on inter-task generalization rather than on intra-task generalization.
@@ VII Further Analysis
(a) After Learning Task 1
(b) After Learning Task 2
(c) After Learning Task 3
(d) After Learning Task 4
(e) After Learning Task 5
Figure 2: T-SNE visualization of memory samples at different learning stages using training set order Yelp $\rightarrow$ AGNews $\rightarrow$ DBpedia $\rightarrow$ Amazon $\rightarrow$ Yahoo. The black-circled data points belong to the last task. Data points with darker colors represent samples from earlier tasks, except (d) after learning Task 4. Task 4 has the same domain as Task 1. Hence, data points with darker colors are belong to the latest task in (d).
@@
Figure 3: Per task accuracy at different learning stages. The dark color indicates high accuracy. From left to right, the color for each task progressively fades away, indicating forgetting happens while learning more tasks. Note that Yelp and Amazon are from the same domain (sentiment analysis).UNC shows a lighter color compared to other AL methods, indicating a higher degree of forgetting. The red-outlined box shows the accuracy on AGNews (Task 2) after learning DBpedia (Task 3).
@@ VII-A Memory Insight
Different learning stages
Fig. 2 presents the T-SNE visualization of memory samples at different learning stages. As show in Figure 2a to 2b, when the proposed model learns a small number of tasks or encounters small label spaces, it focuses on ensuring intra-task generalization. However, as the number of seen tasks increases, the model shifts its focus towards ensuring inter-task generalization as shown in 2d to 2e. Consequently, we observe a phenomenon wherein the memory samples from the last tasks cluster together.
@@ VIII Conclusion
This paper considers a realistic continual learning scenario with limited annotation budgets, namely continual active learning. We devise a novel method to address this problem. Our method dynamically selects worthwhile unlabeled data for annotation to maximize model performance and address continual learning problems. Specifically, it applies MAML with memory replay and consistency regularization to enable efficient learning with limited memory and annotation budgets while mitigating catastrophic forgetting. In experiments, we analyzed the effect of various active learning strategies and memory sample selection methods on model performance. The experimental results shows our method achieves a great performance in a 5-shot continual active learning setup. Furthermore, we demonstrate that random sampling is still the best default strategy for both active learning and memory sample selection to solve 5-shot continual active learning problems. However, we only evaluate our method in a 5-shot continual active learning scenario. In the future work, we can extend our evaluation to a more realistic scenario where the amount of labeled data is {100, 1000, 5000}. In addition, the annotation budget is allocated equally for each task. The annotation budget allocation strategies can be further studied.